Session: Sep 1997 (session 3)
Frequency: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K. Borkowski at Fri Sep 19 11:41:37 METDST 1997) Minor failures
Lost the first 2 minutes and the whole scan at 16:17 UT, both due to problems with the tape drive.

Onsala (John Conway at Fri Sep 19 11:53:56 METDST 1997) Success
No known problems

Medicina (D. Dallacasa at Fri Sep 19 14:48:21 METDST 1997) Success
No known problem

Effelsberg ( at Mon Sep 22 09:08:34 METDST 1997) Success

Noto (C.Trigilio at Mon Sep 22 16:50:04 METDST 1997) Minor failures
One scan (on DA193) lost for antenna cable wrap problem. No other problem found

Westerbork (Tony Foley at Tue Sep 23 17:20:03 METDST 1997) Minor failures
Read headstack sometimes stuck

Cambridge (Alastair Gunn at Tue Sep 23 21:59:51 METDST 1997) Minor failures
Slightly late on first source due to PC crash. Telescope communications failed between 20:28 but telescope may have been on correct source. Possible erroneous tape motion error messages. No other known problems.

Calibration (Fredrik T Rantakyro at Wed Nov 5 15:57:55 MET 1997) Minor failures
The ANTAB formatted TSYS information can be found on VLBEER in the SEP97 directory. The file is called gc018a.antab and contains the latest antenna calibration information.